OpenStudy (anonymous):

What are the x-intercept(s) of the graph of y + 10 = x2 + 3x?

hartnn (hartnn):

to get x-intercepts, just plug in y =0 in y + 10 = x^2 + 3x you will get an easy quadratic equation to solve :)

OpenStudy (uri):

y + 10 = x^2 + 3x y=x2+3x -10 <-Quadratic equation.